The yield of a chemical process is being studied. From a previous experience yield is known to be normally distributed and σ = 3. The past 5 days if plant operation have to resulted in the following percent yields: 91.6,88.75,90.8,89.95, and 91.3. Find a 95% two-sided confidence interval on the true mean yield.

1
Expert's answer
2022-04-28T15:46:03-0400

Sample mean

"\bar{x}=\dfrac{1}{5}(91.6+88.75+90.8+89.95+91.3)=90.48"

 The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(90.48-1.96\times\dfrac{3}{\sqrt{5}}, 90.48+1.96\times\dfrac{3}{\sqrt{5}})"

"=(87.85, 93.11)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "87.85 < \mu < 93.11," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(87.85, 93.11)."